Problem 2 (25 points) SMR For the state of stress at a point, P, as shown. JUNY Plot the state of stress on a Mohrs circle.
Answer #1

30 pa 6x. 8mfа by. - Supa ma tayampa Impa توهكل كملي كتبه 6 to - Galaxy R= Eman S (61-6a2 tengt R=1( 84. 5²+ (302 R = 7.1589